OpenStudy (anonymous):

7 1 ---- + ---- = solve 11w w

OpenStudy (anonymous):

Common Denominator

OpenStudy (anonymous):

7w/11w^2 + 11w/11w^2 = 18w/11w^2 = 18/11w

OpenStudy (anonymous):

\[\frac{7+1(11)}{11w}=\frac{7+11}{11w}\]\[\frac{18}{11w}\]